# Joint Node Selection and Resource Allocation Optimization for Cooperative Sensing with a Shared Wireless Backhaul

### v1

@@ I Introduction
Generally, existing cooperative sensing schemes can be broadly categorized into two types, i.e., information-domain cooperative sensing (IDCS) and signal-domain cooperative sensing (SDCS). In IDCS, each sensing receiver extracts specific location information from the echo signals and then the fusion center (FC) fuses the extracted information for cooperative target localization [10, 11]. The common approaches of IDCS are usually based on the time-of-arrival (TOA) [12, 13], angle-of-arrival (AOA) [14], time-difference-of-arrival (TDOA) [15], received signal strength (RSS) [16], etc. Besides, the localization performance of IDCS is strongly relative to the employed data fusion algorithms [17], which can generally be categorized into two types, i.e., centralized algorithm and distributed algorithm. Centralized algorithms can offer more accurate position estimates in small networks with acceptable complexity [18], while distributed algorithms are also attractive since they can improve the network robustness and scalability [19]. However, the localization accuracy of IDCS is easily affected by non-line-of-sight (NLOS) propagation environments [20]. Besides, according to the law of data processing, extracting location information from echo signals always incurs certain information loss, which may lead to localization performance degradation.
@@
Figure 1: Cooperative sensing system model
@@
Notations: Scalars, vectors and matrices are respectively denoted by lower/upper case, boldface lower case and boldface upper case letters. $\Re(x)$ and $\Im(x)$ denote the real and imaginary parts of a complex number $x$, respectively. $\mathbf{A}^{T}$ and $\mathbf{A}^{-1}$ denote the transpose and inverse of matrix $\mathbf{A}$, respectively. $\mathbf{I}_{n}$ denotes a $n\times n$ identity matrix. $\operatorname{diag}(\mathbf{a})$ denotes a diagonal matrix with the elements in $\mathbf{a}$ being its diagonal elements. $\|\mathbf{a}\|$ denotes the $l_{2}$-norm of vector $\mathbf{a}$, and $\mathbf{A}\succeq\mathbf{B}$ means that $\mathbf{A}-\mathbf{B}$ is a positive semi-definite matrix.
@@ II System Model
where $\omega_{n}(t)\sim\mathcal{CN}(0,\sigma^{2}_{n})$ is the complex Gaussian noise at the $n$-th receiver, $\alpha_{n}=\rho_{n}\xi_{n}$ is the effective reflecting coefficient of the $n$-th receiver, $\rho_{n}$ is the corresponding path-loss coefficient and $\xi_{n}$ is the reflection coefficient, which is assumed to be a complex random variable with unknown (deterministic) amplitude and random phase uniformly distributed between 0 and $2\pi$. In addition, the time delay $\tau_{n}$ corresponding to the $n$-th receiver can be expressed as
@@ III Hybrid Information-Signal Domain Cooperative Sensing
Figure 2: Illustration of the proposed design
@@ III-A ML Estimation of Individual Delay and Effective Reflecting Coefficient
Note that directly solving (5) generally requires a two-dimensional search, which incurs high computational complexity. Thus, we propose to first fix $\tau_{n}$, then the optimal $\hat{\alpha}_{n}$ has a closed-form solution that can be expressed as:
$\hat{\alpha}_{n}=\frac{{\textstyle\sum_{k=1}^{K}}r_{n}(kT_{s})s(kT_{s}-\tau_{n})}{\sqrt{E}{\textstyle\sum_{k=1}^{K}}\left[s(kT_{s}-\tau_{n})\right]^{2}},$ (6)
Substituting (6) back into (5), the ML estimator for the time delay $\hat{\tau}_{n}$ can be obtained as
@@
Since the CRLB offers a theoretical and achievable lower bound of any unbiased estimator, and similar to the model in [28], we assume that the estimated $\hat{\tau}_{n}$ is statistically related to the true parameter $\tau_{n}$ as follows:
@@
where $\omega_{\tau_{n}}\sim\mathcal{N}\left(0,\text{CRLB}_{\hat{\tau}_{n}}\right)$ is a Gaussian noise and its variance $\text{CRLB}_{\hat{\tau}_{n}}$ is the CRLB of $\tau_{n}$ at $\hat{\tau}_{n}$, given by[29]
$\text{CRLB}_{\hat{\tau}_{n}}=\frac{1}{\frac{2E}{\sigma^{2}_{n}}\left|\hat{\alpha}_{n}\right|^{2}\sum_{k=1}^{K}\left[{\frac{\partial s(t)}{\partial t}}\mid_{t=kT_{s}-\hat{\tau}_{n}}\right]^{2}}.$ (9)
@@ III-B Received Signal Quantization at Each Receiver
In this subsection, aiming to reduce the communication overhead (caused by sending all echo signal samples to the FC) and obtain high cooperation gain, we propose to send the received signals sampled around $\hat{\tau}_{n}$ to the FC. Suppose that each receiver has obtained the quantization bit allocation scheme from the FC (the details are provided in Section IV), the KLT based encoding scheme is thus employed to quantize these samples, which is shown to maximize the coding gain for Gaussian sources [30].
@@
Second, in order to efficiently quantize the sampled signals, it is crucial to ultilize the statistical information contained in $\mathbf{r}_{n}$. However, it is very difficult to directly analyze the exact distribution of $\mathbf{r}_{n}$. To make the analysis tractable, we approximate $r_{n}(k_{n}T_{s})$ by considering its first-order Taylor expansion at $\hat{\tau}_{n}$, which can be expressed as
$\begin{split}&r_{n}(k_{n}T_{s})\approx\\ &\sqrt{E}\alpha_{n}\left[\omega_{\tau_{n}}\frac{\partial s(t)}{\partial t}\mid_{t=k_{n}T_{s}-\hat{\tau}_{n}}+s(k_{n}T_{s}-\hat{\tau}_{n})\right]+\omega_{n}(k_{n}T_{s}).\end{split}$ (11)
Note that both $\omega_{n}(k_{n}T_{s})$ and $\omega_{\tau_{n}}$ are Gaussian distributed, we can prove that the approximate value of $\mathbf{r}_{n}$ in (11) also obeys Gaussian distribution with its covariance matrix $\mathbf{Q}_{\mathbf{r}_{n}}$ given by
$\mathbf{Q}_{\mathbf{r}_{n}}=E\cdot\text{CRLB}_{\hat{\tau}_{n}}\mathbf{q}_{\mathbf{r}_{n}}\mathbf{q}_{\mathbf{r}_{n}}^{T}+\frac{1}{2}\sigma^{2}_{n}\mathbf{I}_{2K_{n}},$ (12)
where $\mathbf{q}_{\mathbf{r}_{n}}=[\{\Re(\hat{\alpha}_{n})\frac{\partial s(t)}{\partial t}\mid_{t=k_{n}T_{s}-\hat{\tau}_{n}},k_{n}\in\mathcal{K}_{n}\}$, $\{\Im(\hat{\alpha}_{n})\frac{\partial s(t)}{\partial t}\mid_{t=k_{n}T_{s}-\hat{\tau}_{n}},k_{n}\in\mathcal{K}_{n}\}]^{T}$. The proof is straightforward and is omitted for conciseness.
@@
In the first step, we obtain $\mathbf{Q}_{\mathbf{r}_{n}}=\mathbf{U\Lambda U}^{T}$ via eigenvalue decomposition, where the diagonal matrix $\mathbf{\Lambda}=\textrm{diag}(\gamma_{n1},\gamma_{n2},\cdots,\gamma_{n(2K_{n})})$ collects all the eigenvalues of $\mathbf{Q}_{\mathbf{r}_{n}}$ and that satisfy $\gamma_{n1}=2\gamma_{n2}=\cdots=2\gamma_{n(2K_{n})}=\sigma^{2}_{n}$. Then, by transforming $\mathbf{r}_{n}$ using KLT, we can obtain the following transformed vector:
$\mathbf{r}_{nC}=\mathbf{U}^{T}\mathbf{r}_{n}.$ (13)
Since the covariance matrix of the transformed vector $\mathbf{r}_{nC}$ is the diagonal matrix $\mathbf{\Lambda}$, all the elements in $\mathbf{r}_{nC}$ are independent of each other and satisfy
$\left[\mathbf{r}_{nC}\right]_{j}\sim\mathcal{N}\left(\left[\mathbf{U}^{T}\bar{\mathbf{r}}_{n}\right]_{j},\gamma_{nj}\right),j\in[1,2K_{n}].$ (14)
In the second step, a Lloyd quantizer [31] is applied to quantize $\left[\mathbf{r}_{nC}\right]_{j},j\in[1,2K_{n}]$. Since the locations of the $N$ receivers are different and the impacts of their received signals on sensing performance are distinct, $\left[\mathbf{r}_{nC}\right]_{j},j\in[1,2K_{n}]$ can be quantized using different numbers of quantization bits. Nevertheless, the total number of available quantization bits is constrained by the MAC channel capacity.

### v2

@@ I Introduction
Generally, existing cooperative sensing schemes can be broadly categorized into two types, i.e., information-domain cooperative sensing (IDCS) and signal-domain cooperative sensing (SDCS). In IDCS, each sensing receiver extracts specific location information from the echo signals and then the fusion center (FC) fuses the extracted information for cooperative target localization [10, 11]. The common approaches of IDCS are usually based on the time-of-arrival (TOA) [12, 13], angle-of-arrival (AOA) [14], time-difference-of-arrival (TDOA) [15], received signal strength (RSS) [16], etc. Besides, hybrid IDCS methods which combine TOA, TDOA, AOA, or RSS measurements to improve accuracy and robustness, have gained significant research interests [17, 18]. For example, the work [17] proposed to combine the TOA and AOA information such that both time and angular information measurements are leveraged, which yields a more accurate estimate in complex environments. As for the localization performance of IDCS, it is strongly relative to the employed data fusion algorithms [19], which can generally be categorized into two types, i.e., centralized algorithm and distributed algorithm. Centralized algorithms can offer more accurate position estimates in small networks with acceptable complexity [20], while distributed algorithms are also attractive since they can improve the network robustness and scalability [21]. However, the localization accuracy of IDCS is easily affected by non-line-of-sight (NLOS) propagation environments [22]. Besides, according to the law of data processing, extracting location information from echo signals always incurs certain information loss, which may lead to localization performance degradation.
@@
Figure 1: Cooperative sensing system model.
@@
Notations: Scalars, vectors and matrices are respectively denoted by lower/upper case, boldface lower case and boldface upper case letters. $\mathbb{E}(\mathbf{X})$ denotes the expectation of the random variable $\mathbf{X}$. $\Re(x)$ and $\Im(x)$ denote the real and imaginary parts of a complex number $x$, respectively. $\mathbf{A}^{T}$ and $\mathbf{A}^{-1}$ denote the transpose and inverse of matrix $\mathbf{A}$, respectively. $\mathbf{I}_{n}$ denotes a $n\times n$ identity matrix. $\operatorname{diag}(\mathbf{a})$ denotes a diagonal matrix with the elements in $\mathbf{a}$ being its diagonal elements. $\|\mathbf{a}\|$ denotes the $l_{2}$-norm of vector $\mathbf{a}$, and $\mathbf{A}\succeq\mathbf{B}$ means that $\mathbf{A}-\mathbf{B}$ is a positive semi-definite matrix.
@@ II System Model
where $\omega_{n}(t)\sim\mathcal{CN}(0,\sigma^{2}_{n})$ is the complex Gaussian noise at the $n$-th receiver, $\alpha_{n}=\rho_{n}\xi_{n}$ is the effective reflecting coefficient of the $n$-th receiver, $\rho_{n}$ is the corresponding path-loss coefficient and $\xi_{n}$ is the reflection coefficient, which is assumed to be a complex random variable with unknown (deterministic) amplitude and random phase uniformly distributed between 0 and $2\pi$. We assume that the noises $\omega_{n}(t)$’s at different sensing receivers are statistically independent, due to the spatial separation, varying environmental conditions, and differing hardware factors, etc. This assumption is made to simplify the analysis without loss of generality and is commonly used in similar signal processing frameworks [26, 27]. Besides, the direct signal transmitted from the transmitter to the receivers is not considered in (1) because in many radar and communication systems [28, 29, 23], this direct signal usually arrives first with a different amplitude or phase as compared to the echo signal, thus it can be easily separated or filtered out from the reflected echo signals using techniques such as temporal filtering [28], thanks to the globally known positions of both the transmitter and the receivers. In addition, the time delay $\tau_{n}$ corresponding to the $n$-th receiver can be expressed as
@@ III Hybrid Information-Signal Domain Cooperative Sensing
Figure 2: Illustration of the proposed design.
@@ III-A ML Estimation of Individual Delay and Effective Reflecting Coefficient
Note that directly solving (5) generally requires a two-dimensional search, which incurs high computational complexity. Thus, we propose to first obtain a closed-form solution of $\hat{\alpha}_{n}$ as a function of $\tau_{n}$, given by
$\hat{\alpha}_{n}=\frac{{\textstyle\sum_{k=1}^{K}}r_{n}(kT_{s})s(kT_{s}-\tau_{n})}{\sqrt{E}{\textstyle\sum_{k=1}^{K}}\left|s(kT_{s}-\tau_{n})\right|^{2}}.$ (6)
Then, by substituting (6) back into (5), the ML estimator for the time delay $\hat{\tau}_{n}$ can be obtained as
@@
Finally, once $\hat{\tau}_{n}$ is obtained from (7), it can be substituted back into (6) to compute the corresponding estimate of $\hat{\alpha}_{n}$. The main procedure of the proposed ML algorithm is provided in Fig. 3:
Figure 3: Procedure of the proposed ML algorithm.
The abovementioned approach ensures optimal estimation performance because the solution in equation (6) provides the minimum mean squared error (MMSE) estimate of $\alpha_{n}$ with a fixed $\tau_{n}$, and the maximum likelihood estimation method used in (5) is optimal when there is no prior information about the target. Additionally, this approach simplifies the estimation process by reducing the original two-dimensional search problem to a one-dimensional search problem over $\tau_{n}$.
Since the CRLB offers a theoretical and achievable lower bound of any unbiased estimator, and similar to the model in [33], we assume that the estimated $\hat{\tau}_{n}$ and $\hat{\alpha}_{n}$ are statistically independent and related to the true parameters $\tau_{n}$ and $\alpha_{n}$, respectively, as follows:
@@
$\hat{\alpha}_{n}=\alpha_{n}+\omega_{\alpha_{n}},$ (9)
where $\omega_{\tau_{n}}\sim\mathcal{N}\left(0,\text{CRLB}_{\hat{\tau}_{n}}\right)$ is a Gaussian noise and its variance $\text{CRLB}_{\hat{\tau}_{n}}$ is the CRLB of $\tau_{n}$ shown as follows[34]:
$\text{CRLB}_{\hat{\tau}_{n}}=\frac{1}{\frac{2E}{\sigma^{2}_{n}}\left|\hat{\alpha}_{n}\right|^{2}\sum_{k=1}^{K}\left|{\frac{\partial s(t)}{\partial t}}\mid_{t=kT_{s}-\hat{\tau}_{n}}\right|^{2}}.$ (10)
Similarly, $\omega_{\alpha_{n}}\sim\mathcal{CN}\left(0,\text{CRLB}_{\hat{\alpha}_{n}}\right)$ is a complex Gaussian noise with variance $\text{CRLB}_{\hat{\alpha}_{n}}$, given by
$\text{CRLB}_{\hat{\alpha}_{n}}=\frac{1}{\frac{2E}{\sigma^{2}_{n}}\sum_{k=1}^{K}\left|s(kT_{s}-\hat{\tau}_{n})\right|^{2}},$ (11)
@@ III-B Received Signal Quantization at Each Receiver
In this subsection, aiming to reduce the communication overhead (caused by sending all echo signal samples to the FC) and obtain high cooperation gain, we propose to send the received signals sampled around $\hat{\tau}_{n}$ to the FC. By transmitting the signal samples, we are able to retain the detailed time-domain information that is critical for more precise localization. With the additional time samples as in the proposed method, we are supposed to achieve better resolution and accuracy in localization. Suppose that each receiver has obtained the quantization bit allocation scheme from the FC (the details are provided in Section IV), the KLT based encoding scheme is thus employed to quantize these samples, which is shown to maximize the coding gain for Gaussian sources [35].
@@
Second, in order to efficiently quantize the sampled signals, it is crucial to ultilize the statistical information contained in $\mathbf{r}_{n}$. However, it is very difficult to directly analyze the exact distribution of $\mathbf{r}_{n}$. To make the analysis tractable, we propose to regard $r_{n}(k_{n}T_{s})$ as a binary function of $\tau_{n}$ and $\alpha_{n}$, denoted as $r_{n}(k_{n}T_{s})=g_{n}(\tau_{n},\alpha_{n})$. Then, by considering its first-order Taylor expansion around $\hat{\tau}_{n}$ and $\hat{\alpha}_{n}$, we can approximate $r_{n}(k_{n}T_{s})$ as follows:
$r_{n}(k_{n}T_{s})$ $=g_{n}(\tau_{n},\alpha_{n})$ (13)
$\approx g_{n}(\hat{\tau}_{n},\hat{\alpha}_{n})+(\tau_{n}-\hat{\tau}_{n})\cdot\frac{\partial g_{n}(\hat{\tau}_{n},\hat{\alpha}_{n})}{\partial\tau_{n}}$
$\quad+(\alpha_{n}-\hat{\alpha}_{n})\cdot\frac{\partial g_{n}(\hat{\tau}_{n},\hat{\alpha}_{n})}{\partial\alpha_{n}}$
$=\sqrt{E}\omega_{\tau_{n}}\hat{\alpha}_{n}\frac{\partial s(t)}{\partial t}\mid_{t=k_{n}T_{s}-\hat{\tau}_{n}}$
$\quad+\sqrt{E}(\hat{\alpha}_{n}-\omega_{\alpha_{n}})s(k_{n}T_{s}-\hat{\tau}_{n})+\omega_{n}(k_{n}T_{s}).$
Under the approximation in (13), the approximate value of $\mathbf{r}_{n}$ is obtained as
${\mathbf{r}}_{n}$ $\approx\sqrt{E}\omega_{\tau_{n}}\mathbf{p}_{\mathbf{r}_{n}}-\sqrt{E}\Re(\omega_{\alpha_{n}})\mathbf{q}_{\mathbf{r}_{n}1}-\sqrt{E}\Im(\omega_{\alpha_{n}})\mathbf{q}_{\mathbf{r}_{n}2}$ (14)
$\quad+\sqrt{E}\mathbf{h}_{\mathbf{r}_{n}}+\mathbf{w}_{\mathbf{r}_{n}},$
where
$\mathbf{p}_{\mathbf{r}_{n}}=[\{\Re(\hat{\alpha}_{n}\frac{\partial s(t)}{\partial t}\mid_{t=k_{n}T_{s}-\hat{\tau}_{n}}),k_{n}\in\mathcal{K}_{n}\},$ (15)
$\{\Im(\hat{\alpha}_{n}\frac{\partial s(t)}{\partial t}\mid_{t=k_{n}T_{s}-\hat{\tau}_{n}}),k_{n}\in\mathcal{K}_{n}\}]^{T},$
$\mathbf{q}_{\mathbf{r}_{n}1}=[\{\Re(s(k_{n}T_{s}-\hat{\tau}_{n})),k_{n}\in\mathcal{K}_{n}\},$ (16)
$\{\Im(s(k_{n}T_{s}-\hat{\tau}_{n})),k_{n}\in\mathcal{K}_{n}\}]^{T},$
$\mathbf{q}_{\mathbf{r}_{n}2}=[\{-\Im(s(k_{n}T_{s}-\hat{\tau}_{n})),k_{n}\in\mathcal{K}_{n}\},$ (17)
$\{\Re(s(k_{n}T_{s}-\hat{\tau}_{n})),k_{n}\in\mathcal{K}_{n}\}]^{T},$
$\mathbf{h}_{\mathbf{r}_{n}}=[\{\Re(\hat{\alpha}_{n}s(k_{n}T_{s}-\hat{\tau}_{n})),k_{n}\in\mathcal{K}_{n}\},$ (18)
$\{\Im(\hat{\alpha}_{n}s(k_{n}T_{s}-\hat{\tau}_{n})),k_{n}\in\mathcal{K}_{n}\}]^{T},$
$\mathbf{w}_{\mathbf{r}_{n}}=[\{\Re(\omega_{n}(k_{n}T_{s}),k_{n}\in\mathcal{K}_{n}\},$ (19)
$\{\Im(\omega_{n}(k_{n}T_{s})),k_{n}\in\mathcal{K}_{n}\}]^{T}.$
Since we assume that $\omega_{\alpha_{n}}\sim\mathcal{CN}(0,\textrm{CRLB}_{\hat{\alpha}_{n}})$ is complex Gaussian distributed, its real and imaginary parts are independent and each follows Gaussian distribution $\mathcal{N}(0,\frac{1}{2}\textrm{CRLB}_{\hat{\alpha}_{n}})$. Then, it can be inferred that $\mathbf{r}_{n}$ also obeys Gaussian distribution since $\omega_{n}(k_{n}T_{s})$, $\omega_{\tau_{n}}$, $\Re(\omega_{\alpha_{n}})$ and $\Im(\omega_{\alpha_{n}})$ are Gaussian distributed. Hence, the mean of $\mathbf{r}_{n}$ can be easily obtained as
$\bar{\mathbf{r}}_{n}=\sqrt{E}\mathbf{h}_{\mathbf{r}_{n}},$ (20)
and the covariance matrix $\mathbf{Q}_{\mathbf{r}_{n}}$ of $\mathbf{r}_{n}$ is given by
$\mathbf{Q}_{\mathbf{r}_{n}}$ $=\mathbb{E}[(\mathbf{r}_{n}-\bar{\mathbf{r}}_{n})(\mathbf{r}_{n}-\bar{\mathbf{r}}_{n})^{T}]$ (21)
$=\mathbb{E}(\mathbf{r}_{n}\mathbf{r}_{n}^{T})-\bar{\mathbf{r}}_{n}\bar{\mathbf{r}}_{n}^{T}$
$=E\cdot\mathbf{p}_{\mathbf{r}_{n}}\mathbf{p}_{\mathbf{r}_{n}}^{T}\mathbb{E}(\omega_{\tau_{n}}^{2})+E\cdot\mathbf{q}_{\mathbf{r}_{n}1}\mathbf{q}_{\mathbf{r}_{n}1}^{T}\mathbb{E}(\Re(\omega_{\alpha_{n}})^{2})$
$\quad+E\cdot\mathbf{q}_{\mathbf{r}_{n}2}\mathbf{q}_{\mathbf{r}_{n}2}^{T}\mathbb{E}(\Im(\omega_{\alpha_{n}})^{2})$
$\quad+E\cdot\mathbf{h}_{\mathbf{r}_{n}}\mathbf{h}_{\mathbf{r}_{n}}^{T}+\mathbb{E}(\mathbf{w}_{\mathbf{r}_{n}}\mathbf{w}_{\mathbf{r}_{n}}^{T})-\bar{\mathbf{r}}_{n}\bar{\mathbf{r}}_{n}^{T}$
$=E\cdot\mathrm{CRLB}_{\hat{\tau}_{n}}\mathbf{p}_{\mathbf{r}_{n}}\mathbf{p}_{\mathbf{r}_{n}}^{T}+E\cdot\frac{1}{2}\mathrm{CRLB}_{\hat{\alpha}_{n}}\mathbf{q}_{\mathbf{r}_{n}1}\mathbf{q}_{\mathbf{r}_{n}1}^{T}$
$\quad+E\cdot\frac{1}{2}\mathrm{CRLB}_{\hat{\alpha}_{n}}\mathbf{q}_{\mathbf{r}_{n}2}\mathbf{q}_{\mathbf{r}_{n}2}^{T}+\frac{1}{2}\sigma^{2}_{n}\mathbf{I}_{2K_{n}}.$
@@
In the first step, we obtain $\mathbf{Q}_{\mathbf{r}_{n}}=\mathbf{U}_{n}\mathbf{\Lambda}_{n}\mathbf{U}_{n}^{T}$ via eigenvalue decomposition, where the diagonal matrix $\mathbf{\Lambda}_{n}=\textrm{diag}(\gamma_{n1},\gamma_{n2},\cdots,\gamma_{n(2K_{n})})$ collects all the eigenvalues of $\mathbf{Q}_{\mathbf{r}_{n}}$. Then, by transforming $\mathbf{r}_{n}$ using KLT, we can obtain the following transformed vector:
$\mathbf{r}_{nC}=\mathbf{U}_{n}^{T}\mathbf{r}_{n}.$ (22)
Since the covariance matrix of the transformed vector $\mathbf{r}_{nC}$ is the diagonal matrix $\mathbf{\Lambda}_{n}$, all the elements in $\mathbf{r}_{nC}$ are independent of each other and satisfy
$\left[\mathbf{r}_{nC}\right]_{j}\sim\mathcal{N}\left(\left[\mathbf{U}_{n}^{T}\bar{\mathbf{r}}_{n}\right]_{j},\gamma_{nj}\right),j\in[1,2K_{n}].$ (23)
In the second step, a Lloyd quantizer [36] is applied to quantize $\left[\mathbf{r}_{nC}\right]_{j},j\in[1,2K_{n}]$. Since the locations of the $N$ receivers are different and the impacts of their received signals on sensing performance are distinct, $\left[\mathbf{r}_{nC}\right]_{j},j\in[1,2K_{n}]$ can be quantized using different numbers of quantization bits. Nevertheless, the total number of available quantization bits is constrained by the MAC capacity.